# Multi-Layer Cycle Benchmarking for high-accuracy error characterization

### v1

@@ III Scalable Pauli characterization
The Sparse Pauli-Lindblad (SPL) model, introduced in Ref. [8], provides a scalable framework for characterizing Pauli noise. This model has already been implemented on a 127-qubit device, yielding valuable characterization data that allowed the implementation of noise-aware error mitigation protocols across the entire QPU [15]. Another scalable characterization scheme, known as Cycle Error Reconstruction (CER) [16], has similarly been leveraged to inform error mitigation strategies [20]. Both methods achieve scalability by relying on the reasonable assumption that the most relevant errors have a small, local support, reflecting the QPU’s topology. While our findings are beneficial for both approaches, this work primarly focuses on the SPL model, whose key features are outlined in the remainder of this section.
@@ III.1 The SPL model
The SPL model assumes that $\Lambda_{C}$, the effective Pauli noise map in Eq. (2) associated with the dressed Clifford layer $C$, is generated by a Lindbladian whose jump operators are local and low-weight Pauli operators [8]. Specifically, it can be expressed as
$\Lambda_{C}(\rho)=e^{\mathcal{L}}(\rho)=\exp\sum_{\alpha\in\mathcal{K}}\lambda_{\alpha}^{C}(P_{\alpha}\rho P_{\alpha}-\rho),$ (6)
where the set $\mathcal{K}$ consists of Pauli strings with a support involving up to $w$ neighbouring qubits (with respect to the QPU’s topology). To achieve a full characterization of the error model, all the model coefficients $\lambda^{C}_{\alpha}$ need to be determined. Considering a maximum weight $w=2$ — a common assumption in the literature [8, 15, 29] and which we adopt throughout this paper — the number of model parameters associated with $\Lambda_{C}$ is given by $|\mathcal{K}|=3n+9p$. Here, $n$ is the number of qubits and $p$ is the number of nearest-neighbour pairs. For a square topology, at scale, we have $p\sim 2n$ which leads to the linear scaling $|\mathcal{K}|\sim 21n$.
The model parameters $\lambda_{\alpha}^{C}$ cannot be accessed directly but they are related to the Pauli fidelities, associated with the dressed Clifford layer $C$, via
@@
where $M_{\alpha\beta}=\langle\alpha,\beta\rangle_{sp}$. For brevity, we omitted the superscript $C$ in Eq. (7) as well as in the remainder of the manuscript. It is convenient to generalize Eq. (7) to take into account that, in some cases, only products of Pauli fidelities are known with high accuracy. To this end, we define the full $|\mathcal{K}|$-dimensional vector of model parameters $\vec{\lambda}$  [^3] Implicitly this means that we define an ordering of the set $\mathcal{K}$, which maps a Pauli string $\alpha$ to an integer $i\in\mathbb{N}$ so that the $i$-th element of $\vec{\lambda}$ (denoted by $\lambda_{i}$) is given by $\lambda^{C}_{\alpha}$.. Similarly, we consider an $F$-dimensional vector $\vec{\xi}$ whose $j$-th element consists of a given product of fidelities, associated with the Pauli strings in a given set $\mathcal{X}_{j}$ (which typically corresponds to some orbit), i.e. $\xi_{j}=\prod_{\alpha\in\mathcal{X}_{j}}f_{\alpha}$. Then we can generalize Eq. (7) to
@@
where the logarithm is applied element-wise and $\bar{M}$ is a rectangular $F\times|\mathcal{K}|$ matrix whose $j$-th row reads $\bar{M}_{jk}=\sum_{\alpha\in\mathcal{X}_{j}}M_{\alpha k}$. If enough Pauli fidelities (or products thereof) are known such that the rank of $M$ is equal to $|\mathcal{K}|$, then Eq. (8) can be inverted and the model parameters determined, typically via a non-negative least square fit
$\vec{\lambda}(\vec{\xi})={\rm argmin}_{\lambda_{k}\geq 0}\;||\bar{M}\vec{\lambda}+\log(\vec{\xi})/2||_{2}^{2}.$ (9)
@@
$\Lambda_{C}^{(\beta)}(\rho)=\prod_{k\in\mathcal{K}}\left(w_{k}^{(\beta)}\rho+(1-w_{k}^{(\beta)})P_{k}\rho P_{k}\right)$ (10)
@@
III.2 Effects of limited learnability
@@
Let us consider the characterization of a Clifford layer $C$ composed of parallel CZ gates - a scenario of significant practical relevance [22, 34, 35]. In order to fit the corresponding SPL model, it is sufficient to know [8] (i) all the $3n$ fidelities associated with unit-weight Pauli operators and (ii) all products of fidelity pairs $f_{\alpha}f_{\beta}$ where $\alpha\in\mathcal{K}$ is a weight two Pauli string and $\beta$ is the Pauli string associated to the conjugation of $P_{\alpha}$ under the layer $C$, possibly conjugated with a second Clifford layer, consisting only of 1Q gates  [^5] Indeed, it is required that each one of the two elements of $\beta$ in the support of $\alpha$ is either the identity or equal to the corresponding element of $\alpha$. This can be always ensured by composing the layer $C$ with a second Clifford layer, consisting only of 1Q gates, in the same spirit as the interleaved CB [8]. While the combination of standard and interleaved CB provides us with high-accuracy values for all the fidelity products required in (ii), condition (i) is highly impacted by the limited learnability of Pauli noise and some fidelities will need to be estimated using different, less accurate approaches.
@@
In the previous section, we demonstrated that a single, isolated CZ gate has two unlearnable DOF, which can be conveniently associated with the unit-weight fidelities $f_{IX}$ and $f_{XI}$. Leveraging this fact, along with condition (i), we can immediately compute the number of unlearnable DOF for an SPL characterization of a layer containing $n_{g}$ parallel CZ gates, which totals $2n_{g}$. These can be also conveniently associated with $2n_{g}$ unlearnable unit-weight fidelities, each one of them featuring a single $X$ operator on one qubit belonging to the layer’s support.
Various approaches have been proposed to estimate these unlearnable fidelities. One option involves performing additional experiments using unit-depth ($d=1$) variants of the CB protocols, where the layer to be characterized is applied only once. However, in this case, it is not possible to distinguish the Pauli fidelities of interest from SPAM errors, which can be effectively removed only by assuming perfect state preparation [8, 19, 33]. As a result, the data obtained from this method lacks high accuracy. Another approach assumes symmetries among the Pauli fidelities. For example, it may be assumed that certain unlearnable fidelities are equal to their conjugates under the action of $C$. In the case of a single CZ gate, this corresponds to assuming, for instance, $f_{XI}=f_{XZ}$ and then estimate the individual unlearnable fidelity as $f_{XI}=\sqrt{f_{XI}f_{XZ}}$, where the product on the right hand side is learnable. If this symmetry assumption holds, the SPL model can be fit using only high-accuracy data. However, if the assumption is invalid, this method introduces errors.
@@
Regardless of the chosen strategy for addressing limited learnability, once conditions (i) and (ii) are satisfied, all information about the (products of) fidelities can be encapsulated in the vector $\vec{\xi}$. It is useful to represent $\vec{\xi}$ as a concatenation $\vec{\xi}=(\vec{\Xi},\vec{\zeta})$, where $\vec{\Xi}$ contains only learnable (products of) fidelities, and $\vec{\zeta}$ comprises unlearnable ones. At this stage, the full-rank matrix $\bar{M}$ can be constructed, enabling the SPL model parameters $\vec{\lambda}$ to be determined by inverting Eq. (8). As shown in Refs. [19, 8] and further discussed in Sec.VI, the limited accuracy associated with $\vec{\zeta}$ impacts the overall precision of the scalable error characterization.
@@
So far, we have focused on the characterization of a single layer $C$. However, in realistic scenarios, multiple distinct layers must often be addressed. A particularly relevant example is depicted in Fig. 2, which illustrates a large $n$-qubit QPU with a square topology and four layers of parallel CZ gates, color-coded as blue (B), green (G), red (R), and orange (O). These four layers collectively cover all nearest-neighbor connections on the QPU. Combined with arbitrary 1Q gate layers, they enable the construction of non-trivial circuits that facilitate rapid entanglement distribution across the whole QPU [37]. Such dense configurations are useful for a variety of applications, including simulating the dynamics of lattice models [15] and implementing variational quantum algorithms [38].
@@ III.3 Number of required experiments
Compared to strategies based on depth-one circuits, approaches relying on symmetry assumptions have the significant advantage of not requiring additional experiments. In these cases, the entire experimental runtime is indeed allocated to executing standard and interleaved CB protocols. For the sake of concreteness, we still focus on QPU s with a square lattice topology, as illustrated in Fig. 2. In this scenario, it has been shown that the (products of) fidelities associated with all orbits $\mathcal{R}^{C}_{\alpha}$ for $\alpha\in|\mathcal{K}|$ can be measured using standard CB protocols with just nine distinct state preparation and measurement stages [8]  [^6] This result is not limited to square lattices but applies more generally to topologies where qubits can be ordered such that no qubit is directly connected to more than two of its predecessors.. As for the interleaved CB instances, building on the simple single CZ gate example, it can be shown that just two additional instances of interleaved CB —where extra S gates are added before each CZ gate—are required to capture all the necessary additional (products of) fidelities. Consequently, a single layer of parallel CZ gates can be fully characterized using a total of $11$ CB instances. Here, a single instance refers to a CB experiment with a fixed initial Pauli state and measurement basis. For the four layers shown in Fig. 2, this approach requires a total of $44$ CB instances.
This number can be further reduced to $16$ instances (four per layer) by generalizing the RC procedure to include S gates in the twirling set. This extension symmetrizes the $X$ and $Y$ components of the noise, reducing the amount of information that needs to be extracted and thereby decreasing the experimental runtime [40]. It is important to emphasize that the discussion here focuses solely on the number of unique CB instances required—an essential aspect of implementing the procedure efficiently [40]. However, this does not address the sampling complexity of the characterization strategies, which is also an important factor in determining the overall runtime. A more comprehensive discussion of the sampling complexity associated with Pauli noise characterization can be found in Ref. [41].
@@ IV Multi-layer CB
In realistic scenarios, quantum circuits typically involve more than a single (dressed) Clifford layer, requiring the characterization of an entire set $\mathcal{C}$ of distinct layers [15, 8]. The conventional approach to this problem involves characterizing each unique layer individually, following the procedure outlined earlier. In this section, we propose augmenting the conventional approach with the novel MLCB strategy that leverages circuits retaining the fundamental CB structure but relying on building blocks that consist of combinations of different layers, picked from the set $\mathcal{C}$. A simple-yet-illustrative example is shown in Fig. 3(a), where a combination of two layers, depicted in blue and green, is repeated $d$ times. In the remainder of this Section, we demonstrate the potential of this technique by focusing at first on the characterization of a simple system and then discussing its generalization to more complicated and relevant scenarios.
Figure 3: (a) Illustration of the MLCB protocol. The structure is analogous to the one shown in Fig. 1 with the important difference that, here, the “building block” which is repeated $d$ times is not a single Clifford layer but a combination of more layers (in this specific case two, shown in blue and green). (b) Simple scenario where MLCB allows to measure with high accuracy the product $f^{B}_{IXI}f^{G}_{XZX}$ of two fidelities, belonging to two different layers, which cannot be accessed with conventional CB protocols. (c) Simple system consisting of three qubits and two layers.
@@ IV.1 Simple example
The simple setup shown in Fig. 3(c) consists of three qubits arranged in a line and numbered from 00 to $2$. The goal is to characterize the two depicted layers, which correspond to a single CZ between qubits 00 and $1$ (blue layer) and between qubits $1$ and $2$ (green layer). The conventional approach, described in the previous Section, prescribes to focus on each layer individually and perform CB protocols to determine with high accuracy all the (products of) fidelities related to the orbits that involve the Pauli strings $\alpha\in\mathcal{K}$. Since each layer consists of a single CZ gate, there are four unlearnable DOF, which can be identified with the set of weight-one fidelities
@@
Here we use the notation $f^{L}_{\alpha}$ for the Pauli fidelities, where $\alpha$ is a Pauli string (ordered according to the increasing qubit indexes) and $L\in\{B,G\}$ indicates the associated layer, i.e. the blue (B) one or the green (G) one. To complete the characterization task, all the fidelities in (11) must be estimated using low-accuracy methods, such as symmetry assumptions or unit-depth variants of CB protocols.
Remarkably, the MLCB experiment shown in Fig. 3(b) enhances the characterization process by enabling the high-accuracy characterization of one additional DOF. To demonstrate this, it is crucial to determine the behavior of a Pauli operator $P_{\alpha}$ under conjugation by the alternating sequence of blue and green layers, which constitutes the core of MLCB circuits. The first $m\geq 0$ layers of this sequence, starting from the leftmost blue layer, are associated with the unitary
$\bar{U}^{(m)}=\begin{cases}(U_{B}U_{G})^{m/2}&\text{ if }m\text{ is even },\\ U_{B}(U_{G}U_{B})^{(m-1)/2}&\text{ if }m\text{ is odd },\end{cases}$ (12)
where $U_{B}$ and $U_{G}$ are the unitary associated with the blue and green layer, respectively. Let us define the sequence of Pauli strings $\alpha^{(m)}$ by requiring that
$P_{\alpha^{(m)}}=\bar{U}^{(m)}P_{\alpha}[\bar{U}^{(m)}]^{\dagger}.$ (13)
In the specific example of Fig. 3(b), with the initial string $\alpha=\alpha^{(0)}=XIX$, the sequence is $2$-periodic: $\alpha^{(2k)}=XIX$ and $\alpha^{(2k+1)}=XZX$ for $k\geq 0$. If we prepare a $+1$ eigenstate of operator $P_{XIX}$, implement $d$ pairs of the blue and green layer and then measure (with readout twirling) the expectation value $\tilde{\mathsf{O}}_{d}^{XIX}$ of the same operator, we obtain
$\tilde{\mathsf{O}}_{d}^{XIX}=A_{XIX}(f^{B}_{XIX}f^{G}_{XZX})^{d}.$ (14)
In analogy with the standard CB protocols, the contribution of SPAM errors is encapsulated in the prefactor $A_{XIX}$ and a simple exponential fit of data obtained for different $d$ provide us with an high-accuracy estimation of the product $o_{3}=f^{B}_{XIX}f^{G}_{XZX}$.
The two fidelities in $o_{3}$ differ form the ones in Eq. 11 and, if we were considering a general Pauli noise model, they would therefore not help in imposing constraints on the fidelities in Eq. 11. Importantly, however, in the context of effective noise models, which feature a limited number of parameters and thus allow for scalable characterization, not all the Pauli fidelities are independent quantities. Specifically, considering the SPL model, we show in Appendix B that
$f^{B}_{XII}=\frac{f^{B}_{XIX}}{f^{B}_{IIX}}$ (15)
and
$\begin{split}f^{G}_{IIX}=\frac{f^{G}_{XZI}(f^{G}_{IZX}f^{G}_{IIX})}{f^{G}_{XZX}f^{G}_{IZI}},\end{split}$ (16)
which leads to
$\mu_{1}^{BG}\equiv\frac{f^{B}_{XII}}{f^{G}_{IIX}}=o_{3}\,\frac{f^{G}_{IZI}}{f^{B}_{IIX}f^{G}_{XZI}(f^{G}_{IZX}f^{G}_{IIX})}.$ (17)
The meaning of the notation $\mu_{1}^{BG}$ will become apparent later. All the terms on the RHS can be determined with high accuracy, using MLCB for $o_{3}$ and standard CB protocols for all the other terms. Consequently, MLCB grants us access to an additional DOF, which can be identified with the fidelity ratio $\mu_{1}^{BG}$, which was unlearnable under the conventional approach alone.
To extend this result to more complex scenarios, it is crucial to formalize the concept of equivalence between functions of fidelities. Specifically, we define two functions as equivalent if one can be fully determined from the other using only high-accuracy data, i.e. (products of) learnable fidelities. Within this framework, Eq. (17) demonstrates that $\mu_{1}^{BG}$ and $o_{3}$ are equivalent, providing a foundation for applying this principle in more intricate cases.

### v2

@@ III Scalable Pauli characterization
Estimating all the $4^{n}-1$ parameters of the Pauli noise map associated with the execution of the layer $C$ on a $n$-qubit system is not possible at scale. It is therefore necessary to rely on simpler, effective Pauli models featuring a number of parameters that scales more favourably with $n$. A natural and established way to achieve this is to leverage the locality of the physical noise processes and the resulting sparsity of the model parameters. A prominent example here is the marginalization procedure, a key component of the Cycle Error Reconstruction (CER) scheme [7] which has been successfully used to inform experimental realizations of noise-aware error mitigation technique [15]. Another well-known scalable framework for characterizing Pauli noise is the so-called Sparse Pauli-Lindblad (SPL) model, introduced in Ref. [45] and successfully used to perform noise-aware error mitigation on large QPUs [32, 18].
@@ III.1 The SPL model
The limited learnability of Pauli noise directly also impacts effective and scalable models [46, 18, 7]. Consequently, the implementation of MLCB is beneficial, irrespective of the specific Pauli model under consideration. For the sake of concreteness, however, the remainder of this paper focuses primarily on the SPL model, which we introduce briefly below.
The SPL model assumes that $\Lambda_{C}$ is generated by a Lindbladian whose jump operators are local and low-weight Pauli operators [46], that is $\Lambda_{C}=e^{\mathcal{L}}$ with
$\mathcal{L}(\rho)=\sum_{\alpha\in\mathcal{K}}\lambda_{\alpha}^{C}(P_{\alpha}\rho P_{\alpha}-\rho).$ (4)
Here the set $\mathcal{K}$ consists of all Pauli strings with a support involving up to $w$ neighbouring qubits (with respect to the QPU’s topology). This also allows the model to capture crosstalk terms. Considering a maximum weight $w=2$ — a common assumption in the literature [46, 32, 17] and which we adopt throughout this paper — the number of model parameters associated with $\Lambda_{C}$ is given by $|\mathcal{K}|=3n+9p$, regardless of the number of gates in the layer, with $n$ the number of qubits and $p$ the number of nearest-neighbour pairs. For a square topology, at scale, we have $p\sim 2n$ which leads to the linear scaling $|\mathcal{K}|\sim 21n$.
The model parameters $\lambda_{\alpha}$ cannot be accessed directly but they are related to the Pauli eigenvalues via
@@
where $M_{\alpha\beta}=\langle\alpha,\beta\rangle_{sp}$ and we omitted the superscript $C$ for brevity. Because of limited learnability, sometimes only products of Pauli eigenvalues are known with high accuracy. It is thus convenient to define the full $|\mathcal{K}|$-dimensional vector of model parameters $\vec{\lambda}$  [^1] Implicitly this means that we define an ordering of the set $\mathcal{K}$, which maps a Pauli string $\alpha$ to an integer $i\in\mathbb{N}$ so that the $i$-th element of $\vec{\lambda}$ (denoted by $\lambda_{i}$) is given by $\lambda^{C}_{\alpha}$.. Similarly, we consider an $F$-dimensional vector $\vec{\xi}$ whose $j$-th element consists of a given product of eigenvalues, associated with the Pauli strings in a given set $\mathcal{X}_{j}$ (which typically corresponds to some orbit), i.e. $\xi_{j}=\prod_{\alpha\in\mathcal{X}_{j}}f_{\alpha}$. Then we can generalize Eq. (5) to
@@
where the logarithm is applied element-wise and $\bar{M}$ is a rectangular $F\times|\mathcal{K}|$ matrix whose $j$-th row reads $\bar{M}_{jk}=\sum_{\alpha\in\mathcal{X}_{j}}M_{\alpha k}$. If enough Pauli eigenvalues (or products thereof) are known such that the rank of $M$ is equal to $|\mathcal{K}|$, then Eq. (6) can be inverted and the model parameters determined, typically via a non-negative least square fit
[…]
@@
$\Lambda_{C}^{(\beta)}(\rho)=\prod_{k\in\mathcal{K}}\left(w_{k}^{(\beta)}\cdot+(1-w_{k}^{(\beta)})P_{k}\cdot P_{k}\right)\rho$ (8)
@@
III.2 Effects of limited learnability on the SPL model
@@
Let us consider the characterization of a Clifford layer $C$ composed of parallel CZgates — a scenario of significant practical relevance [1, 2, 36]. In order to fit the corresponding SPL model, it is sufficient to know [46] (i) all the $3n$ eigenvalues associated with weight-one Pauli operators and (ii) all products of pairs of $f_{\alpha}f_{\beta}$ where $\alpha\in\mathcal{K}$ is a weight-two Pauli string and $P_{\beta}$ is the conjugation of $P_{\alpha}$ under the layer $C$ and, possibly, additional single-qubit Clifford gates  [^3] Indeed, it is required that each one of the two elements of $\beta$ in the support of $\alpha$ is either the identity or equal to the corresponding element of $\alpha$. This can be always ensured by composing the layer $C$ with a second Clifford layer, consisting only of single-qubit gates [46]. While the standard CB protocol provides us with high-accuracy values for all the products of eigenvalues required in (ii), condition (i) is highly impacted by the limited learnability of Pauli noise and some eigenvalues will need to be estimated using alternative, low-accuracy approaches.
@@
In the previous section, we demonstrated that a single, isolated CZ gate has two unlearnable DOF, which can be conveniently associated with the weight-one Pauli eigenvalues $f_{IX}$ and $f_{XI}$. Leveraging this fact, along with condition (i), we can immediately compute the number of unlearnable DOF for an SPL characterization of a layer containing $n_{g}$ parallel CZgates, which totals $2n_{g}$. These can also be conveniently associated with $2n_{g}$ unlearnable weight-one eigenvalues, each one of them associated with a single $X$ operator on one qubit belonging to the layer’s support.
Several methods have been proposed to estimate these unlearnable quantities. One approach is the non-SPAM-robust $d=1$ CB variant mentioned earlier. Another common—but theoretically unjustified—strategy relies on assuming symmetries among the Pauli eigenvalues. For instance, in the case of a single CZ gate, one might assume that unlearnable eigenvalues are equal to their conjugates under the action of $C$, such as $f_{XI}=f_{XZ}$. This allows the estimation of individual unlearnable eigenvalues via $f_{XI}=\sqrt{f_{XI}f_{XZ}}$, where the product on the right-hand side is learnable. However, since this assumption lacks rigorous justification, the method can introduce significant errors and yields low accuracy.
@@
Regardless of the chosen strategy for addressing limited learnability, once conditions (i) and (ii) are satisfied, all information about the (products of) eigenvalues can be encapsulated in the vector $\vec{\xi}$. It is useful to represent $\vec{\xi}$ as a concatenation $\vec{\xi}=(\vec{\Xi},\vec{\zeta})$, where $\vec{\Xi}$ contains only learnable (products of) eigenvalues, and $\vec{\zeta}$ comprises unlearnable ones. At this stage, the full-rank matrix $\bar{M}$ can be constructed, enabling the SPL model parameters $\vec{\lambda}$ to be determined by inverting Eq. (6). As shown in Refs. [9, 46] and further discussed in Sec.VI, the limited accuracy associated with $\vec{\zeta}$ impacts the overall precision of the scalable error characterization.
@@
So far, we have focused on the characterization of a single layer $C$. However, in realistic scenarios, multiple distinct layers must often be addressed. A particularly relevant example is depicted in Fig. 2, which illustrates a large $n$-qubit QPU with a square topology and four layers of parallel CZgates, color-coded as blue (B), green (G), red (R), and orange (O). These four layers collectively cover all nearest-neighbor connections on the QPU. Combined with arbitrary single-qubit gate layers, they enable the construction of non-trivial circuits that facilitate rapid entanglement distribution across the whole QPU [3]. Such dense configurations are useful for a variety of applications, including simulating the dynamics of lattice models [32] and implementing variational quantum algorithms [39].
@@ III.3 Number of required experiments
Compared to strategies based on depth-one circuits, approaches relying on symmetry assumptions have the significant advantage of not requiring additional experiments. In these cases, the entire experimental runtime is indeed allocated to executing the standard CB protocol. For the sake of concreteness, we still focus on QPU s with a square lattice topology, as illustrated in Fig. 2. In this scenario, it has been shown that the (products of) eigenvalues associated with all orbits $\mathcal{R}^{C}_{\alpha}$ for $\alpha\in\mathcal{K}$ can be measured using the standard CB protocol with just eleven CB instances, a single instance referring to given state preparation and measurement stages and potential interleaving of extra single-qubit Clifford gates [46]  [^4] This result is not limited to square lattices but applies more generally to topologies where qubits can be ordered such that no qubit is directly connected to more than two of its predecessors.. . For the four layers shown in Fig. 2, this approach requires a total of $44$ CB instances.
This number can be further reduced to $16$ instances (four per layer) by generalizing the PT procedure to include S gates in the twirling set. This extension symmetrizes the $X$ and $Y$ components of the noise, reducing the amount of information that needs to be extracted and thereby decreasing the experimental runtime [47]. It is important to emphasize that the discussion here focuses solely on the number of unique CB instances required—an essential aspect of implementing the procedure efficiently [47]. However, this does not address the sampling complexity of the characterization strategies, which is also an important factor in determining the overall runtime. A more comprehensive discussion of the sampling complexity associated with Pauli noise characterization can be found in Ref. [10].
@@ IV Multi-layer CB
In realistic scenarios, quantum circuits typically involve more than a single (dressed) Clifford layer, requiring the characterization of an entire set $\mathcal{C}$ of distinct layers. The conventional approach to this problem involves characterizing each unique layer individually, following the procedure outlined earlier [32, 46]. In this section, we propose augmenting the conventional approach with the novel MLCB strategy that leverages circuits retaining the fundamental CB structure but relying on building blocks that consist of combinations of different layers, picked from the set $\mathcal{C}$. A simple yet illustrative example is shown in Fig. 3(a), where a combination of two layers, depicted in blue and green, is repeated $d$ times. Note that each layer is independently twirled, as shown by the white boxes indicating single-qubit twirling gates.
To properly describe the MLCB protocol, it is convenient to introduce the definition of a multi-layer orbit associated with the pair of layers $L_{1},L_{2}\in\mathcal{C}$. It generalizes Eq. (3) and reads
$\mathcal{R}_{\alpha}^{L_{1}L_{2}}=\left(P_{\alpha^{(0)}},P_{\alpha^{(1)}},\dots,P_{\alpha^{(m-1)}}\right)$ (9)
where
$P_{\alpha^{(j)}}=\phi(\bar{U}_{L_{1}L_{2}}^{(j)})[P_{\alpha}]$ (10)
and the unitary
[…]
captures the effect of two alternating layers $L_{1},L_{2}\in\mathcal{C}$, associated with unitaries $U_{L_{1}}$ and $U_{L_{2}}$. The cardinality $m\geq 2$ of the multi-layer orbit $\mathcal{R}_{\alpha}^{BG}$ is the smallest even integer such that $\phi(\bar{U}_{L_{1}L_{2}}^{(m)})[P_{\alpha}]=P_{\alpha}$. Analogously to the standard CB protocol, by preparing a +1+1 eigenstate of the operator $P_{\alpha}$ and measuring its expectation value after the implementation of $m/2$ pairs of twirled layers, the MLCB protocol can thus measure in a SPAM-robust way and with multiplicative precision the product of the Pauli eigenvalues
$\mathcal{F}_{\alpha}^{L_{1}L_{2}}=\prod_{j=0}^{m/2-1}f^{B}_{\alpha^{(2j)}}f^{G}_{\alpha^{(2j+1)}},$ (12)
where the Pauli strings are defined according to Eq. (10).
In the remainder of this Section, we demonstrate how this technique can be used to boost the learnability of Pauli noise models by focusing at first on the characterization of a simple system and then discussing its generalization to more complicated and relevant scenarios.
Figure 3: MLCB protocol: (a) The structure is analogous to the one shown in Fig. 1 with the important difference that, here, the “building block” which is repeated $d$ times is not a single Clifford layer but a combination of more layers (in this specific case two, shown in blue and green). Single-qubit gates used for twirling are indicated in white. (b) Simple scenario where MLCB allows one to measure with high accuracy the product $f^{B}_{XIX}f^{G}_{XZX}$ of two eigenvalues, belonging to two different layers, which cannot be accessed with conventional CB protocols. For the sake of simplicity, single-qubit gates used for twirling are now explicitly shown. (c) Simple system consisting of three qubits and two layers.
@@ IV.1 Simple example
The simple setup shown in Fig. 3(c) consists of three qubits arranged in an open chain and numbered from 0 to $2$. The goal is to characterize the two depicted layers, which correspond to a single CZbetween qubits 0 and $1$ (blue layer) and between qubits $1$ and $2$ (green layer). The conventional approach, described in the previous Section, prescribes to focus on each layer individually and perform CB protocols to determine with high accuracy all the (products of) eigenvalues related to the orbits that involve the Pauli strings $\alpha\in\mathcal{K}$. Since each layer consists of a single CZ gate, there are four unlearnable DOF, which can be identified with the set of weight-one eigenvalues
@@
Here we use the notation $f^{L}_{\alpha}$ for the Pauli eigenvalues, where $\alpha$ is a Pauli string (ordered according to the increasing qubit indexes) and $L\in\{B,G\}$ indicates the associated layer, i.e. the blue (B) one or the green (G) one. To complete the characterization task, all the eigenvalues in (13) must be estimated using low-accuracy methods, such as symmetry assumptions or unit-depth variants of CB protocols.
Notably, the MLCB experiment shown in Fig. 3(b) significantly advances the characterization process by enabling the high-accuracy determination of an additional DOF. This is accomplished by analyzing the orbit $\mathcal{R}_{XIX}^{BG}=(P_{XIX},P_{XZX})$, which yields a precise estimate of the product of individually unlearnable Pauli eigenvalues:
$\mathcal{F}^{BG}_{XIX}=f^{B}_{XIX}f^{G}_{XZX}.$ (14)
These eigenvalues correspond to high-weight, non-local Pauli operators, which would be unrelated to the unlearnable DOF in Eq.(13) in the context of a general Pauli noise model. Within the SPL framework, however, not all Pauli eigenvalues are independent from each other and, as we demonstrate in Appendix C, the knowledge of $\mathcal{F}^{BG}_{XIX}$ provided by MLCB is equivalent to the knowing the ratio $\mu_{1}^{BG}\equiv{f^{B}_{XII}}/{f^{G}_{IIX}}$ between two unlearnable weight-one eigenvalues in Eq. (13). Specifically, this relationship is expressed as
$\mu_{1}^{BG}\equiv\frac{f^{B}_{XII}}{f^{G}_{IIX}}=\mathcal{F}^{BG}_{XIX}\,\frac{(f^{G}_{IZI})}{(f^{B}_{IIX})(f^{G}_{XZI})(f^{G}_{IZX}f^{G}_{IIX})}$ (15)
where all the terms in round brackets are learnable and can be determined with high-accuracy using conventional CB protocols. The non-trivial structure of Eq. (15) arises because the SPL model accounts for crosstalk, which introduces correlated errors also between neighboring qubits 0 and $1$.
Building on this insight, we generalize the approach, presenting a systematic method to identify relevant multi-layer orbits. This allows MLCB to impose high-accuracy constraints on ratios of otherwise unlearnable weight-one Pauli eigenvalues for arbitrary combinations of Clifford layers.